how do we relate the angles in radians and a real no.?

Grade:11

2 Answers

Ankit Jaiswal
7 years ago
Hi ayush,
the main thing to know about is that 
\piradian = 180\degree
and hence 
1\degree = \pi/180 radian and hence x\degree = x*\pi/180 radian
also
1 radian =180/ \pi\degree and x radian =x*180/ \pi\degree
now the value of \pi \approx 3.14 or 22/7 
on substituting the value of pi i the above equation u can change degree and radians into real no.
I hope this helps 
thanks
Ankit Jaiswal
7 years ago
 
Hi ayush
this question is prety easy 
\piradian = 180\degree
and hence 
1\degree = \pi/180 radian and hence x\degree = x*\pi/180 radian
also
1 radian =180/ \pi\degree and x radian =x*180/ \pi\degree
now the value of \pi \approx 3.14 or 22/7 
on substituting the value of pi i the above equation u can change degree and radians into real no.
I hope this helps 
thanks